jim and bob are panning for gold in leadville colorado. ratio of gold found by jim to gold found by bob is 4:7. if the two men together find a total of 55 pieces of gold. find how many pieces of gold each man found. use the 5-step method and check your work
Answer by checkley77(12844)

4n+7n=11n number of shares.
11n=55
n=55/11
n=5 pieces of gold per share.
4*5=20 pieces for Jim.
7*5=35 pieces for Bob.
Proof:
20+35=55
55=55